# Generalization Beyond Data Imbalance: A Controlled Study on CLIP for Transferable Insights

### v2

@@ 1 Introduction
The development of contrastive language-image pre-training (CLIP) [65, 35, 90, 42, 54] has demonstrated unprecedented success in learning generalizable representations, empowering zero-shot vision tasks and robustness to natural distributional shifts. This success can be primarily attributed to the effective use of large-scale uncurated image captioning datasets collected from the web. A recent trend involves delving into the distribution of these datasets and explicitly introducing interventions to the curation process to create better data for training [28, 88]. However, limited research has been conducted on analyzing the distribution of concepts/classes in these datasets and the behavior of CLIP under varying distributions. This work thus starts by presenting a concept-centric analysis of existing web-scale image-text datasets and models pre-trained accordingly (Footnote 2).
@@
Figure 1: Per-class statistics of image-text datasets and models trained on top. (a) A highly imbalanced class distribution is shared across datasets. [^2] MetaCLIP [88] is relatively more balanced than other datasets due to concept re-balancing in curation.(b) Compared to supervised learning ( SL), CLIP’s performance (measured by $\bullet$ accuracy) is less biased by data frequency, and the classifier is notably uncorrelated (measured by model’s number of $\bullet$ prediction per class). Besides, the correlation narrows as data scales up. Both aspects indicate implicit re-balancing mechanisms exist in CLIP.
@@ 2 Related work
3 What makes CLIP more robust to pre-training data imbalance?
@@ 5 Limitations, future work, and broader impacts
Future work. Our findings cover insights in language supervision, pretext task, data scaling, and concept scaling, but only a small portion are validated in application. One direction for future work is to explore the use of language supervision and open-world data in recognition models. Besides, a recent work [41] finds Adam optimizer to outperform (stochastic) gradient descent on heavy-tailed data, which can be another factor in CLIP’s robustness and is worth further exploration. On the other hand, we are interested in extending our discovery to the open-world detection and segmentation tasks to see if our findings still hold under these more challenging scenarios. Furthermore, as we have analyzed in our study, language supervision plays an important role in achieving such robustness to the data imbalance, thus we are also interested in studying whether or not similar traces of generalization exist in (multi-modal) large language models (e.g., Llama [80], BLIP-2 [43], LLaVA [45], etc.) and potentially reveal the reason for their effectiveness in future studies.
@@ Acknowledgments
A Controlled Study on CLIP for Transferable Insights
Supplementary Material
(a) Correlation statistics of models pre-trained on ImageNet-Captions.
(b) Correlation statistics of models pre-trained on LAIONet.
(c) Correlation statistics of models pre-trained on YFCC-15M.
Figure 11: Which is a better indicator for per-class statistics? (a) For less imbalanced IN-Caps, both Pearson’s $r$ [63] and Spearman’s $\rho$ [75] can model the correlation between statistics well. (b & c) For extremely imbalanced datasets (e.g., LAIONet, YFCC-15M, and other web datasets), Peason’s $r$ may fail even if class frequencies are processed to log scale. In contrast, Spearman’s $\rho$ remains robust.
@@ A.5 Is data imbalance not a concern for CLIP?
Figure 15: CLIP still can be biased by pre-training data. It is relatively more robust than SL (a), and the bias reduces to some extent as data scales (b vs. a), but the tail classes still underperform.

### v3

# What Makes CLIP More Robust to Long-Tailed Pre-Training Data? A Controlled Study for Transferable Insights

@@ 1 Introduction
The development of contrastive language-image pre-training (CLIP) [68, 36, 93, 44, 57] has demonstrated unprecedented success in learning generalizable representations, empowering zero-shot vision tasks and robustness to natural distributional shifts. This success can be primarily attributed to the effective use of large-scale uncurated image captioning datasets collected from the web. A recent trend involves delving into the distribution of these datasets and explicitly introducing interventions to the curation process to create better data for training [29, 91]. However, limited research has been conducted on analyzing the distribution of concepts/classes in these datasets and the behavior of CLIP under varying distributions. This work thus starts by presenting a concept-centric analysis of existing web-scale image-text datasets and models pre-trained accordingly (Footnote §).
@@
Figure 1: Per-class statistics of image-text datasets and models trained on top. (a) A highly imbalanced class distribution is shared across datasets.§§footnotemark: §(b) Compared to supervised learning ( SL), CLIP’s performance (measured by $\bullet$ accuracy) is less biased by data frequency, and the classifier is notably uncorrelated (measured by model’s number of $\bullet$ prediction per class). Besides, the correlation narrows as data scales up. Both aspects indicate implicit re-balancing mechanisms exist in CLIP.
@@
11footnotetext: MetaCLIP [91] is relatively more balanced than other datasets due to concept re-balancing in curation.
@@ 2 Related work
3 What makes CLIP more robust to long-tailed pre-training data?
@@ 5 Limitations, future work, and broader impacts
Future work. Our findings cover insights in language supervision, pretext task, data scaling, and concept scaling, but only a small portion are validated in application. One direction for future work is to explore the use of language supervision and open-world data in recognition models. Besides, a recent work [43] finds Adam optimizer to outperform (stochastic) gradient descent on heavy-tailed data, which can be another factor in CLIP’s robustness and is worth further exploration. On the other hand, we are interested in extending our discovery to the open-world detection and segmentation tasks to see if our findings still hold under these more challenging scenarios.
Furthermore, as we have analyzed in our study, language supervision plays an important role in achieving such robustness to the data imbalance, thus we are also interested in studying whether or not similar traces of generalization exist in (multi-modal) large language models (e.g., Llama [83], BLIP-2 [45], LLaVA [48], etc.). However, despite being trained on large-scale data with language supervision, recent works show that LLM/MLLMs still suffer from long-tailed training data [46, 37], and their performance is highly correlated with the frequency that corresponding knowledge appeared in training [1, 95]. This indicates that generative models might be intrinsically more prone to long-tailed data than contrastive models like CLIP, and injecting rebalancing mechanisms into the generative process could be valuable for future explorations.
@@ A.5 Is data imbalance not a concern for CLIP?
Figure 15: CLIP can still be biased by pre-training data. It is relatively more robust than SL (a), and the bias reduces to some extent as data scales (b vs. a), but the tail classes still underperform.